Pre-Laboratory Assignment 1. Why shouldnt you look directly at burning magnesium? 2. A compound composed of N and O has 2.5 moles of O for every mole of N. Write the empirical formula for this compound. 3. A student used the Procedure of this experiment to determine the empirical formula for an oxide of tin (Sn). He collected the following data: 22.36 mass of crucible, cover, and Sn, g mass of crucible, cover, and tin oxide,g 24.67 24.98 (a) Calculate the mass of Sn in the tin oxide product. (b) Calculate the mass of O in the tin oxide.
Answer #1

Answer 1. The burning magnesium ribbon produces light of sufficient intensity to cause temporary loss of sight. The burning ofmagnesium in air produces intense heat which can cause burns and initiate combustion in flammable materials and can be dangerous to our eyes too.

Answer 2.

1 mole of N = 2.5 mole of O

2 mole of N = 5 mole of O

Therefore, compound formula = N2O5.

Answer 3.

a.) Mass of Tin = W2-W1 = 24.67 - 22.36 = 2.31 grams.

b.) Mass of O = Mass of Tin Oxide - Mass of Tin = (W3-W1)-W2 = (24.98-22.36)-2.31 = 0.31 grams.

where W1 = mass of crucible and cover

W2 = mass of crucible, cover and Sn

W3 = mass of crucible, cover and tin oxide